# Privacy Perspectives and Practices of Chinese Smart Home Product Teams

### v1

@@ 4.3.2 Suggested social interventions
Many participants stressed the need for collaboration between smart home companies, government agencies, academic institutions, and privacy experts to create effective strategies and standards aimed at improving user privacy awareness and protection. For example, P10-PM highlighted the value of academic involvement in connecting industry practices with research insights, stating: “Universities and companies could hold annual meetings to discuss and integrate academic insights into industry practices.” However, participants admitted that merging academic expertise with industry practices can be challenging, as P15-EN noted: “Certain experts don’t fully grasp the practical realities of the industry, which makes the standards they develop feel helpless in practice. It often feels like “non-experts” are driving the development.”
In addition, some participants emphasized the importance of better privacy education for the general public in China, observing that many Chinese users tend to view privacy rights as less important for ordinary people than for celebrities or the wealthy. Several participants recommended that the government launch community outreach initiatives to raise privacy awareness and encourage best practices among users. For example, P14-EN suggested: “If possible, local sub-district communities could hold privacy protection workshops that require the participation of all community members.” P6-UX additionally highlighted the importance of SOEs in leading privacy education initiatives: “Major SOEs should educate the public on privacy, data usage, and protection strategies. We could include privacy handbooks alongside smart home device manuals and distribute these materials to device owners and purchasers.” Further, P10-PM emphasized the need to integrate privacy education into the Nine-Year Compulsory Education curriculum888 Compulsory Education Law of the People’s Republic of China, stating that doing so would foster awareness of privacy from a young age and equip future generations with the knowledge to better understand and safeguard their privacy rights.
@@ 4.3.3 Suggested legal interventions
Many participants highlighted the need for clearer legal regulations around data management. For example, P10-PM pointed out that vague legal terminology made it difficult to determine when privacy breaches occurred. P1-EN suggested: “Clearly defined legislation with stipulated penalties for privacy violations could prompt companies to adopt a more cautious approach during development.” In contrast, P5-UX, P12-PM, and P19-LE argued that laws should be intentionally broad to encompass a wider array of innovative practices, ensuring they stayed effective as new data collection methods emerged. A small number of participants shared their experiences designing smart homes for overseas customers. For example, P14-EN noted: “Privacy laws in China are not as comprehensive as GDPR. We should work on adapting international standards to fit the local context. Although Chinese privacy regulations have been revised to keep pace with technological and social changes, they often fall behind, resulting in gaps and challenges in enforcement.” P1-EN, P2-PM, and P27-MS talked about China’s strict regulatory environment, which restricted foreign services and closely monitored the privacy practices of domestic companies, to protect national security. Additionally, most participants highlighted the central government’s crucial role in overseeing data privacy and security, emphasizing its significant impact on how companies handled privacy. As P9-UX remarked, “Organizations at the community level, such as ’juweihui’, are essential to raise awareness about privacy and security, helping individuals to better grasp applicable regulations.” Some participants highlighted the crucial role national organizations could play in protecting user data by establishing privacy standards. For example, P1-EN suggested: “Certain national standards bodies could establish a multilevel system for privacy data management, classifying devices into low-, medium-, or high-risk groups, with each category subject to distinct compliance and verification requirements.”
Summary. Participants used technical measures to comply with Chinese privacy laws/regulations, but acknowledged a shortage of adequate social and legal interventions, leading some to suggest improvements in these areas.
@@ 5 Discussion
5.1 Main Findings
Participants showed a strong commitment to legal compliance and ethical design in the early design and development stages, regularly holding review meetings to ensure compliance with Chinese privacy and data protection laws, treating this as part of privacy-related requirements gathering. However, in the final testing phases, usability and functionality were prioritized over privacy. Many participants, especially UX designers often had little to no formal training in legal compliance, suggesting limited understanding of privacy laws and regulations, whereas most product managers displayed greater knowledge of legal compliance and data practices. Some participants’ personal experiences with smart home devices helped them develop empathy for users. However, their practices were heavily shaped by Chinese cultural values, which tend to minimize individual privacy in favor of convenience and collective benefits (RQ1).
Our participants encountered three main challenges: 1) low public awareness of privacy issues affecting at-risk groups, 2) finding a balance between privacy and user experience, and 3) managing power dynamics within the Chinese multi-user smart home ecosystem. These challenges were especially significant in multi-user settings, where coordinating consent and privacy among multiple stakeholders was complex. We also found that Chinese cultural norms, shaped by Confucian values such as filial piety, limited children’s ability to contest parental monitoring through smart home devices out of respect for parental authority. Furthermore, some companies granted emergency police access to all user registration and facial data, and government or public security agencies could forcibly obtain this information, potentially circumventing company-imposed restrictions (RQ2).
@@ 5.2 Contributions over Previous Studies
While some of our findings align with existing research showing that software developers often prioritize functionality and usability over security and privacy [89, 6, 121, 115, 119, 29, 117, 114, 5, 44, 66], we specifically examine these challenges in the context of smart home design and development. This setting introduces distinct privacy concerns due to its intimate environment and complex multi-user interactions. In addition, our participants, who held various product team roles beyond software development, navigated additional complexities such as balancing diverse stakeholder needs, managing hierarchical family relationships, and adhering to the evolving and complex regulatory framework in a non-WEIRD country (see §4.2.2). For example, whereas Western research typically suggests individually tailored privacy controls to safeguard each household member’s autonomy, our findings show participants created paternalistic features allowing parents to remotely monitor their children through ‘family supervision devices’ (e.g., child desk lamps). This design approach reflects Chinese cultural values such as filial piety and collective well-being, differing from Western norms where privacy preferences are usually managed independently by each household member without a central family authority controlling surveillance.
@@
Compared to prior studies on smart home design and development outside China (e.g., [24, 25, 12, 10, 11, 12]) emphasizing challenges such as limited resources, regulatory restrictions, and cultural attitudes, our research uniquely examines how Chinese smart home product teams navigate their distinct regulatory and socio-cultural environment (see §4.1 and §4.2). Our findings not only shed light on the Chinese context but also offer insights applicable to other non-WEIRD regions facing similar cultural and regulatory dynamics. For example, in Southeast and South Asia, culturally adaptive privacy-by-design frameworks can help align product development with local norms and evolving standards [46, 97] (see §5.4.5). Also, we enrich the literature by providing in-depth empirical evidence about the privacy perspectives/practices of product team members in diverse roles and organizations, including start-ups and state-owned enterprises, offering a comprehensive view of the smart home industry. Although earlier Chinese privacy research has focused on user perspectives [58, 80], our study distinctively highlights the viewpoints of product team members, illustrating how they manage privacy within their cultural context (see §4.1.2, §4.2.2).
@@
Furthermore, China’s extensive government oversight of data access imposes compliance burdens, often forcing product teams to provide data to state authorities [76, 125]. Our study builds on this prior work by providing a comprehensive perspective on how product teams manage these constraints and trade-offs in their day-to-day design and development processes. Unlike Western contexts—where companies must adhere to strict privacy regulations like GDPR [73]—Chinese small and medium-sized enterprises typically rely on legal teams’ ecosystem-level checklists rather than adopting systematic privacy engineering or threat modeling practices [35, 116]; see §4.2.2. Although Western software developers often balance privacy with usability, the widespread acceptance of state surveillance in China reduces the emphasis on comprehensive privacy threat modeling, highlighting a culturally-specific norm remaining underexplored in Western-centric research.
Finally, our participants prioritized compliance with Chinese privacy laws, in contrast to prior research showing Western developers often fail to meet GDPR requirements [14]. However, unlike GDPR, Chinese laws emphasize national security over individual privacy.
@@ 5.3 Chinese Social, Cultural, and Legal Factors
5.3.1 Influence by Confucianism
Participants believed that many Chinese users were unconcerned about privacy, shaped by socio-cultural norms (§4.1.2). Some felt ordinary people’s privacy mattered less than the wealthy’s—a view rooted in Confucian values prioritizing hierarchy and social roles, where privacy is often seen as less critical for lower-status individuals [56, 57, 97, 123, 131].
While the notion of “yin si” (privacy) is less central in Chinese culture compared to Western liberal thought, Confucian values—especially those emphasizing family duties—influenced the decisions of smart home product team members. Teams reflected these values by designing features that promoted family cohesion and enabled parental or elder care (see §4.2.2). For example, they added functions allowing children to remotely monitor and support their aging parents, aligning with filial piety norms and parental respect [54]. Devices also included features for parental monitoring—such as tracking study progress. Participants noted that users generally viewed such monitoring as a helpful parenting tool, not as invasive. This aligns with Confucian values that tie academic achievement to filial duty [87, 133, 79, 104], and where respect for hierarchy may discourage youth from resisting such monitoring practices [124]. These choices contrast with Western design norms, which prioritize individual privacy and autonomy, including stricter protections for children’s data and requirements like explicit consent, reflecting a stronger focus on individual rights over collective family control.
@@ 5.3.3 Community-based education
To raise privacy awareness, we highlight the importance of community-based agencies and sub-districts as grassroots platforms (see §4.3.3). Backed by Residents Committees (RCs, juweihui), which play a key role in community management and education, these platforms can reach hundreds to thousands of residents [129]. We recommend broad community training programs and privacy protection lectures for all community members; and distributing privacy-focused materials such as handbooks and posters. Although primarily aimed at the general public, product team members—being part of the broader population—can also benefit from these resources. Furthermore, product teams would benefit from formal company-led training or workshops (see §5.4.4).
@@ 5.3.4 Clear and enforceable privacy laws
Participants (see §4.1.1, §4.2.2, §4.3.3) noted that applying foreign legal standards in the context of Chinese smart home product design is challenging due to vague privacy regulations, unclear data handling protocols, and the difficulty of proving privacy breaches in court [33]. These gaps leave users vulnerable, with little legal recourse and high burdens of proof. Although China’s PIPL outlines clearer definitions for concepts like consent and data portability [137], its enforcement is largely left to individual companies. This has led to inconsistent compliance practices, with firms often relying on internal interpretations and informal guidance rather than standardized protocols. Government-corporate power dynamics—particularly regarding state access to user data—further complicate efforts to balance regulatory compliance, business responsibilities, and individual privacy. To respond, participants advocated for clearer legal definitions and practical implementation frameworks aligned with China’s regulatory context and industry conditions (e.g., government-approved third-party enforcement). A tiered compliance structure—tailored to company size, sector, and risk level—would offer a pragmatic path toward enforceable and trusted privacy protections in the Chinese smart home industry.
@@ 5.4.1 Implement clear privacy guidelines
Participants highlighted difficulties in embedding strong privacy and security features into smart home devices, driven by rising user expectations for transparent data practices in China (see §4.2.1 and §4.1.2), echoing trends observed in WEIRD contexts [14, 37, 24, 25]. Although national standards like GB/T 35273-2020 prohibit default privacy policy consent checkboxes [90], many UX designers in our study struggled with implementing appropriate data handling and avoiding privacy-invasive elements like dark patterns [21, 42], making it harder to communicate privacy terms clearly. To address this, we recommend creating detailed, practical design guidelines that incorporate privacy-by-design principles [23, 55, 100] and emphasize privacy from the early stages of development, especially in multi-user environments. These resources should feature real-world use cases and expert input (e.g., privacy champions [115]) to support designing with privacy as a core principle, fostering a culture where privacy is built-in rather than added later.
@@ 5.4.2 Enhance privacy compliance in small businesses
We found that most large companies—including SOEs, PEs, and MNCs—generally followed privacy regulations during product development (see §4.1.1, §4.2.2, and §4.3.3), restricting team access to sensitive user data and enforcing internal privacy policies (see §4.1.2). Large SOEs, under stricter government scrutiny and public accountability, adhered to more rigorous standards. While large PEs showed more flexibility, they still complied with national laws. In contrast, smaller businesses often prioritized business needs over privacy and legal compliance due to limited resources, echoing findings in prior research [65, 120, 12]. To support smaller companies, we suggest that Chinese regulators develop clear and enforceable frameworks aligned with domestic privacy laws and national standards (e.g., PIPL – Chapter 6). These should include specific guidance for small businesses, clarify enforcement authority, offer support resources, propose strategies for managing limited capacities, and introduce incentives for compliance. Establishing community-based agencies and industry associations can further promote adherence to legal requirements.
@@ 5.4.3 Consider power dynamics and at-risk groups
Based on findings from §4.1.1 and §4.2.2, we observed limited awareness among product teams regarding the privacy needs of at-risk groups in China. Men in Chinese households frequently held decision-making power and device ownership in smart homes, positioning women in the same household as ‘passenger user’ [68, 41, 110, 36, 10], echoing broader patriarchal norms. Teams also struggled to engage with at-risk populations like domestic workers [19, 107, 9, 51] due to limited access, ethical complexities, and workers’ reluctance to share information [98]. These barriers hinder the creation of inclusive smart home technologies. To address this, we recommend ethical, user-centered design guidelines that reflect the privacy needs of all household members, especially women and vulnerable users. Government-led efforts (e.g., by the CAC and Women’s Federation) and local bodies like RCs can offer targeted support [50]. We also call on researchers to build frameworks that systematize and translate at-risk users’ concerns into actionable insights for product design.

### v2

@@ 4.3.2 Suggested social interventions
Many participants stressed the need for collaboration between smart home companies, government agencies, academic institutions, and privacy experts to create effective strategies and standards aimed at improving user privacy awareness and protection. For example, P10-PM highlighted the value of academic involvement in connecting industry practices with research insights, stating: “Universities and companies could hold annual meetings to discuss and integrate academic insights into industry practices.” However, participants admitted that merging academic expertise with industry practices can be challenging, as P15-EN noted: “Certain experts don’t fully grasp the practical realities of industry, which makes the standards they develop feel useless in practice. It often feels like “non-experts” are driving the development.”
In addition, some participants emphasized the importance of better privacy education for the general public in China, believing that many Chinese users tend to view privacy rights as less important for ordinary people than for celebrities or the wealthy. Several participants recommended that the government launch community outreach initiatives to raise privacy awareness and encourage best practices among users. For example, P14-EN suggested: “If possible, local sub-district communities could hold privacy protection workshops that require the participation of all community members.” P6-UX additionally highlighted the importance of SOEs in leading privacy education initiatives: “Major SOEs should educate the public on privacy, data usage, and protection strategies. We could include privacy handbooks alongside smart home device manuals and distribute these materials to device owners and purchasers.” Further, P10-PM emphasized the need to integrate privacy education into the Nine-Year Compulsory Education curriculum999 Compulsory Education Law of the People’s Republic of China, stating that doing so would foster awareness of privacy from a young age and equip future generations with the knowledge to better understand and safeguard their privacy rights.
@@ 4.3.3 Suggested legal interventions
Many participants highlighted the need for clearer legal regulations around data management. For example, P10-PM pointed out that vague legal terminology made it difficult to determine when privacy breaches occurred. P1-EN suggested: “Clearly defined legislation with stipulated penalties for privacy violations could prompt companies to adopt a more cautious approach during development.” In contrast, P5-UX, P12-PM, and P19-LE argued that laws should be intentionally broad to encompass a wider array of innovative practices, ensuring they stayed effective as new data collection methods emerged. A small number of participants shared their experiences designing smart homes for overseas customers. For example, P14-EN noted: “Privacy laws in China are not as comprehensive as GDPR. We should work on adapting international standards to fit the local context. Although Chinese privacy regulations have been revised to keep pace with technological and social changes, they often fall behind, resulting in gaps and challenges in enforcement.” P1-EN, P2-PM, and P27-MS talked about China’s strict regulatory environment, which restricted foreign services and closely monitored the privacy practices of domestic companies, to protect national security.
Additionally, most participants highlighted the central government’s crucial role in overseeing data privacy and security, emphasizing its significant impact on how companies handled privacy. As P9-UX remarked, “Organizations at the community level, such as ’juweihui’, are essential to raise awareness about privacy and security, helping individuals to better grasp applicable regulations.” Some participants highlighted the crucial role national organizations could play in protecting user data by establishing privacy standards. For example, P1-EN suggested: “Certain national standards bodies could establish a multi-level system for data management, classifying devices into low-, medium-, or high-risk categories, with each category subject to distinct compliance and verification requirements.”
Summary. Participants used technical measures to comply with Chinese privacy laws and regulations, but acknowledged a shortage of adequate social and legal interventions, leading some to suggest improvements in these areas.
@@ 5 Discussion
5.1 Key Findings
Participants showed a strong commitment to legal compliance and ethical design in the early design and development stages, regularly holding review meetings to ensure compliance with Chinese privacy and data protection laws, treating this as part of privacy-related requirements gathering. However, in the final testing phases, usability and functionality were prioritized over privacy. Privacy–convenience trade-offs also emerged when teams collected data that was not essential to core device functions (e.g., for personalization), where UX was sometimes prioritized over privacy. Many participants, especially UX designers often had little to no formal training in legal compliance, suggesting limited understanding of privacy laws and regulations, whereas most product managers displayed greater knowledge of legal compliance and data practices. Some participants’ personal experiences with smart home devices helped them develop empathy for users. However, their practices were heavily shaped by Chinese cultural values, which tend to minimize individual privacy in favor of convenience and collective benefits (RQ1).
Our participants encountered three main challenges: 1) low public awareness of privacy issues affecting at-risk groups, 2) finding a balance between privacy and UX, and 3) managing power dynamics within the Chinese multi-user smart home ecosystem. These challenges were especially significant in multi-user settings, where coordinating consent and privacy among multiple stakeholders was complex. We also found that Chinese cultural norms, shaped by Confucian values such as filial piety, limited children’s ability to contest parental monitoring through smart home devices out of respect for parental authority. Furthermore, some companies granted emergency police access to all user registration and facial data, and government or public security agencies could forcibly obtain this information, potentially circumventing company-imposed restrictions (RQ2).
@@ 5.2 Contributions
Although some of our findings align with existing research showing that software developers often prioritize functionality and usability over security and privacy [78, 5, 106, 100, 104, 29, 102, 99, 4, 42, 59], we specifically examine the challenges faced by product teams in the context of smart home design and development. This setting introduces distinct privacy concerns due to its intimate environment and complex multi-user interactions. In addition, our participants—who held diverse product team roles beyond software development—navigated additional complexities, including balancing the needs of different stakeholder groups, managing hierarchical family relationships, and operating within an evolving and often ambiguous regulatory framework in a non-WEIRD context. In some cases, their practices—shaped by local cultural and familial norms—extended beyond complying with legal requirements, underscoring how regulatory frameworks, cultural values, and contextual factors jointly influenced team decision-making (§4.2.2). For example, whereas Western research typically suggests individually tailored privacy controls to safeguard each household member’s autonomy, our findings show participants created paternalistic features allowing parents to remotely monitor their children through ‘family supervision devices’ (e.g., child desk lamps). This design approach reflects Chinese cultural values such as filial piety and collective well-being, differing from Western norms where privacy preferences are usually managed independently by each household member without a central family authority controlling surveillance. Although our study does not provide a direct comparison between the privacy practices of Western and non-Western smart home product teams, it highlights culturally-embedded design choices within Chinese smart home teams that may contrast with dominant assumptions in Western-centric privacy literature.
@@
Compared to prior studies on smart home product design and development outside China (e.g., [25, 26, 9, 10]), our research uniquely examines how Chinese smart home product teams navigate their distinct regulatory and socio-cultural environment (§4.1 and §4.2). Whereas Albayaydh and Flechais [11] emphasized localization and contextual norms, our study shifts the focus inward to company practices, analyzing Chinese product teams and tracing how privacy is operationalized through requirements gathering, design trade-offs, and the prioritization of compliance with Chinese privacy laws. Moreover, our findings do not only shed light on the Chinese context but also offer insights relevant to other non-WEIRD regions facing similar cultural and regulatory dynamics. For example, in Southeast and South Asia, culturally adaptive privacy-by-design frameworks can help align product development with local norms and evolving standards [44, 84]5.4.5). Also, we enrich the literature by providing in-depth empirical evidence about the privacy perspectives/practices of product team members in diverse roles and organizations, including start-ups and state-owned enterprises, offering a comprehensive view of the smart home industry. Although earlier Chinese privacy research has focused on end-user privacy perspectives [53, 71], our study distinctively highlights the points of view of product team members, illustrating how they manage privacy within their cultural context (§4.1.2, §4.2.2).
@@
Furthermore, China’s extensive government oversight of data access imposes significant compliance burdens, often compelling product teams to provide data to state authorities [67, 111]. Building on this prior work, our study offers a comprehensive perspective on how product teams navigate these constraints and trade-offs in their daily design and development practices. By distinguishing between essential and non-essential data, we show how Chinese product teams’ decisions produced different types of trade-offs: essential data often raised questions of minimization and proportionality, whereas non-essential data heightened tensions between added convenience and user privacy. We also found that Chinese small and medium-sized companies typically relied on legal teams’ ecosystem-level checklists, rather than adopting systematic privacy engineering or threat modeling practices [34, 101] (see §4.2.2). The widespread acceptance of state surveillance in China reduces the emphasis on comprehensive privacy threat modeling, highlighting a culturally specific norm that remains underexplored in Western-centric research.
@@ 5.3 Chinese Social, Cultural, and Legal Factors
5.3.1 Influence by Collectivism and Confucianism
Collectivist values could, in principle, motivate a shared responsibility among product team members to protect the privacy of end-users (as well as bystanders) [22]. Yet, most of our participants did not frame their roles in these terms; instead, they typically aligned their privacy practices with legal compliance or market demands. Thus, our findings do not necessarily suggest that privacy is unimportant in China, or that Chinese smart home device end-users prioritize national security over individual privacy; rather, they illustrate how product teams’ perceptions of Chinese end-users, laws, and cultural norms shape the way privacy is operationalized in practice. Further, some participants felt that ordinary people’s privacy mattered less than that of the wealthy—a view rooted in Confucian values, prioritizing hierarchy and social roles, where privacy is often considered less critical for lower-status individuals [51, 52, 84, 109, 116].
While the notion of “yin si” (privacy) is less central in Chinese culture compared to Western liberal thought, Confucian values—especially those emphasizing family duties—influenced the decisions of smart home product team members. Teams reflected these values by designing features that promoted family cohesion and enabled parental or elder care (§4.2.2). For example, they added functions allowing children to remotely monitor and support their aging parents, aligning with filial piety norms and parental respect [49]. Devices also included features for parental monitoring, such as tracking study progress. Participants noted that users generally viewed such monitoring as a helpful parenting tool, not as invasive. This aligns with Confucian values that tie academic achievement to filial duty [76, 118, 70, 90], and where respect for hierarchy may discourage youth from resisting such monitoring practices [110]. These choices contrast with Western design norms, which prioritize individual privacy and autonomy, including stricter protections for children’s data and requirements like explicit consent, reflecting a stronger focus on individual rights over collective family control.
@@ 5.3.3 Community-based education
To raise privacy awareness, we highlight the importance of community-based agencies and sub-districts as grassroots platforms (§4.3.3). Backed by Residents Committees (RCs, juweihui), which play a key role in community management and education, these platforms can reach hundreds to thousands of residents [115]. We recommend broad community training programs and privacy protection lectures for all community members; and distributing privacy-focused materials such as handbooks and posters. Although primarily aimed at the general public, product team members, being part of the broader population, can also benefit from these resources. Furthermore, product teams would benefit from formal company-led training or workshops (§5.4.4).
@@ 5.3.4 Clear and enforceable privacy laws
Participants (§4.1.1, §4.2.2, §4.3.3) noted that applying foreign legal standards in the context of Chinese smart home product design is challenging due to vague privacy regulations, unclear data handling protocols, and the difficulty of proving privacy breaches in court [33]. These gaps leave users vulnerable, with little legal recourse and high burdens of proof. Although China’s PIPL outlines clearer definitions for concepts like consent and data portability [123], its enforcement is largely left to individual companies. This has led to inconsistent compliance practices, with firms often relying on internal interpretations and informal guidance rather than standardized protocols. Government-corporate power dynamics, particularly regarding state access to user data, further complicate efforts to balance regulatory compliance, business responsibilities, and individual privacy. To respond, participants advocated for clearer legal definitions and practical implementation frameworks aligned with China’s regulatory context and industry conditions (e.g., government-approved third-party enforcement). A tiered compliance structure, customized to company size, sector, and risk level, would offer a pragmatic path toward enforceable and trusted privacy protections in the Chinese smart home industry.
@@ 5.4.1 Implement clear privacy guidelines
Participants highlighted difficulties in embedding strong privacy and security features into smart home devices, driven by rising user expectations for transparent data practices in China (§4.2.1 and §4.1.2), echoing trends observed in WEIRD contexts [13, 36, 25, 26]. Although national standards such as GB/T 35273-2020 prohibit default privacy policy consent checkboxes [79], many UX designers in our study struggled with implementing appropriate data handling and avoiding privacy-invasive elements like dark patterns [19, 40], making it harder to communicate privacy terms clearly. To address this, we recommend creating detailed, practical design guidelines that incorporate privacy-by-design principles [24, 50, 87] and emphasize privacy from the early stages of development, especially in multi-user environments. These resources should feature real-world use cases and expert input (e.g., privacy champions [100]) to support designing with privacy as a core principle, fostering a culture where privacy is built-in rather than added later.
@@ 5.4.2 Enhance privacy compliance in small businesses
We found that most large companies—including SOEs, PEs, and MNCs—generally followed privacy regulations during product development (§4.1.1, §4.2.2, and §4.3.3), restricting team access to sensitive user data and enforcing internal privacy policies (§4.1.2). Large SOEs, under stricter government scrutiny and public accountability, adhered to more rigorous standards. While large PEs showed more flexibility, they still complied with national laws. In contrast, smaller businesses often prioritized business needs over privacy and legal compliance due to limited resources, echoing findings in prior research [58, 105, 11]. To support smaller companies, we suggest that Chinese regulators develop clear and enforceable frameworks aligned with domestic privacy laws and national standards (e.g., PIPL – Chapter 6). These should include specific guidance for small businesses, clarify enforcement authority, offer support resources, propose strategies for managing limited capacities, and introduce incentives for compliance. Establishing community-based agencies and industry associations can further promote adherence to legal requirements.
@@ 5.4.3 Consider power dynamics and at-risk groups
Based on findings from §4.1.1 and §4.2.2, we observed limited awareness among product teams regarding the privacy needs of at-risk groups in China. Men in Chinese households frequently held decision-making power and device ownership in smart homes, positioning women in the same household as ‘passenger user’ [61, 39, 96, 35, 9], echoing broader patriarchal norms. Teams also struggled to engage with at-risk populations like domestic workers [17, 93, 8, 47] due to limited access, ethical complexities, and workers’ reluctance to share information [85]. These barriers hinder the creation of inclusive smart home technologies. To address this, we recommend ethical, user-centered design guidelines that reflect the privacy needs of all household members, especially women and vulnerable users. Government-led efforts (e.g., CAC and Women’s Federation) and local bodies such as RCs can offer targeted support [46]. We also call on researchers to build frameworks that systematize and translate at-risk users’ concerns into actionable insights for product design.